OpenStudy (anonymous):

Juliet calculated her age to be eleven and five twelfths years. What number is equivalent to eleven and five twelfths? eleven point four one six eleven point four one six with a bar over four one six eleven point four one six with a bar over the six eleven point four one six with a bar over one six

OpenStudy (anonymous):

11+(5/12) = 11 + (0.416 repeating) = 11.0416